OpenStudy (anonymous):

The hypotenuse of a right triangle is 17. The lengths of the two legs are whole numbers. What are they?

OpenStudy (dn_20):

The length that have 17 as hypotenuse is 8 and 15 \[8^2++ 15^2 = 17^2\] \[64 +225 = 289\]